OpenStudy (anonymous):

A tin can (right circular cylinder) with top and bottom is to have volume V . What dimensions (the radius of the bottom and the height) give the minimum total surfacearea?

OpenStudy (anonymous):

I get \[r=\sqrt[3]{V/(2 \pi)}\], that sound right?

OpenStudy (xapproachesinfinity):

this is calculus question correct

OpenStudy (anonymous):

correct

OpenStudy (xapproachesinfinity):

we are minimizing \[A=2\pi r^2+2\pi r h\]

OpenStudy (xapproachesinfinity):

actually your question is vague somewhat lol

OpenStudy (anonymous):

yes, so I used the folume formula to solve for h, and substituted into the formula above to put in terms of r only, took that derivative with respect to r, and came up with the above

OpenStudy (anonymous):

so, I understand V to be a constant

OpenStudy (xapproachesinfinity):

you are looking for two dimensions that minimize the surfacearea not just r

OpenStudy (anonymous):

yes, and they will be in terms of V

OpenStudy (anonymous):

\[V=h(\pi)r^2 , A=h2(\pi)r + 2(\pi)r^2\]

OpenStudy (xapproachesinfinity):

well write h in terms of v and r so you have only one variable for the A

OpenStudy (anonymous):

\[A=2V/r+2 \pi r^2\]

OpenStudy (xapproachesinfinity):

do the derivative

OpenStudy (xapproachesinfinity):

solve for A'=0

OpenStudy (anonymous):

\[r=\sqrt[3]{V/(2 \pi)}\]

OpenStudy (xapproachesinfinity):

alright if that's the answer should be good

OpenStudy (anonymous):

which equation would I best use to solve to find h

OpenStudy (xapproachesinfinity):

the Volume equation

OpenStudy (anonymous):

any chance you'd walk me through your steps in doing so?

OpenStudy (xapproachesinfinity):

once r is found h is found easily that's what that equation is doing

OpenStudy (xapproachesinfinity):

you already did my friend you don't have to do anything else

OpenStudy (anonymous):

what do you get for h?

OpenStudy (xapproachesinfinity):

you just use v=pir^2h to find h so you have two equation one for r and one for h

OpenStudy (xapproachesinfinity):

i didn't do it you just use \[V=\pi r^2h \Longrightarrow r=\sqrt{V/\pi h}\] you replace r with this

OpenStudy (xapproachesinfinity):

and solve for h

OpenStudy (anonymous):

r is the third root of RHS

OpenStudy (xapproachesinfinity):

\[\sqrt{V/\pi h}=\sqrt[3]{V/2\pi} \Longrightarrow V/h\pi=(V/2\pi)^{2/3 }\]

OpenStudy (xapproachesinfinity):

continue with this

OpenStudy (xapproachesinfinity):

the goal is to find two formulas one for r and one for h and that is all

OpenStudy (anonymous):

thanks for the help X

OpenStudy (xapproachesinfinity):

no problem :)